Leonardo, who is married but files separately, earns $60,000 of taxable income. He also has $6,500 in city of Tulsa bonds. His wife, Theresa, earns $33,000 of taxable income. If Leonardo earned an additional $37,000 of taxable income this year, what would be the marginal tax rate on the extra income for 2019? (Use tax rate schedules)
Answer #1
Tax on $60000 9058.5 =4543+(60000-39475)*22%
Taxable income 97000 =60000+37000
Tax on $97000 17454.5 =14382.5+(97000-84200)*24%
Marginal tax rate 22.69% =(17454.5-9058.5)/37000